# Universal Mott quantum criticality in a modified periodic Anderson model

### v1

@@ III.3.2 DC conductivity and resistivity scaling
where $\rho_{c}^{\text{DC}}(T)$ is the resistivity measured precisely at the critical point and $f$ is a universal scaling function. The characteristic energy scale, $T_{0}$, depends on the deviation from the critical interaction, $\delta U=U-U_{c}$, via the power law $T_{0}\sim|\delta U|^{\nu z}$, with $\nu z$ being the dynamical critical exponent. The SBHM exhibits such scaling with an exponent $\nu z=0.56\pm 0.01$[23] when calculated using IPT and CTQMC impurity solvers within DMFT. However, subsequent studies on the same model have reported slightly different values: DMFT+IPT calculations by Vučičević et al.[25] yield $\nu z=0.6$, while DMFT+NRG results by Eisenlohr et al.[24] report $\nu z=0.66\pm 0.1$. This variation in the reported exponent is understood to arise from different methodological choices for the Widom line, the trajectory in the temperature-interaction plane used as a reference for the scaling analysis. In this paper, we choose a straight Widom line, similar to the approach discussed in Eisenlohr et al[24].
Since our model undergoes a continuous Mott transition, we expect a similar quantum critical scaling in the electrical resistivity. Here, we perform the scaling analysis at a fixed inter-orbital coupling $t_{\perp}$, while varying the Hubbard interaction $U$. Fig. 9 shows the scaled resistivity $\rho^{\text{DC}}(T)/\rho^{\text{DC}}_{c}(T)$ as a function of the scaled temperature $T/T_{0}$ for different values of $U$. Here we choose a straight line, $U^{*}(T)$ above the QCP as a Widom line[24]. The scaling temperature $T_{0}$ exhibits a power-law dependence, characterized by an exponent of approximately $\nu z_{\text{met}}=0.76$ on the metallic side and $\nu z_{\text{ins}}=0.66$ on the insulating side, as shown in the inset of Fig. 9. The disparity in the exponents across the two sides may originate from the choice of the Widom line, as discussed by Eisenlohr et al.[24], who reported $\nu z\in(0.5,0.8)$ for different Widom line definitions in the SBHM. Our results for the MPAM fall within this range, indicating that the critical scaling of the MPAM is consistent with that of the SBHM.
This finding suggests that the MPAM and SBHM may belong to the same universality class. Consistent with this interpretation, experimental studies on systems exhibiting either first-order or continuous Mott transitions report critical exponents within these bounds. Therefore, the MPAM can be regarded as a prototypical model for exploring genuine Mott quantum criticality.
@@ IV CONCLUSIONS AND OUTLOOK
In this paper, using NRG, we have demonstrated that a modified periodic Anderson model exhibits Mott quantum criticality, characterized by quantum critical scaling in resistivity with a critical exponent comparable to that of the SBHM. This suggests that MPAM and SBHM may belong to the same universality class. Importantly, the critical temperature $T_{c}$ of the MPAM is zero. We have also shown that the MPAM displays $\omega/T$ scaling, a hallmark of local quantum criticality. Furthermore, we identified isosbestic points in the optical conductivity and showed that the isosbestic points vanish upon approaching the QCP. Analytical expressions for the MIR peak and the charge excitation peak in the optical conductivity were derived in terms of model parameters. We observed that the MIR peak reaches its minimum at the QCP.
A continuous Mott-like metal–insulator transition has been observed in the Falicov–Kimball model [71, 72], although its critical exponent $\nu z$ is approximately twice that found in the SBHM. Several attempts have been made to tune the finite-temperature critical point of the SBHM to $T=0$ using disorder [73] or particle–hole asymmetry, but these have been unsuccessful. In contrast, the MPAM features a continuous, symmetry-unbroken transition at zero temperature from a Fermi liquid to a Mott insulating state, while showing a comparable critical exponent $\nu z$ as the SBHM. These findings open up several avenues for future exploration. A central question is whether our results define the universality class for Mott quantum criticality. Another key direction is to explore if a suitable parameter can tune the SBHM to a zero-temperature transition, thereby realizing the physics of the MPAM. Furthermore, investigating the role of disorder in this quantum critical system remains a compelling challenge for future study.
@@ Acknowledgements.
N. S. V. and S. K. K. acknowledge funding from the National Supercomputing Mission and JNCASR. S. K. K. acknowledges financial support from the DST-INSPIRE fellowship.
@@ Appendix A Analytic expressions for roots of the Green’s function
$G_{ff}$ $(\omega,\epsilon)=\frac{1}{\omega-\Sigma_{\text{f}}(\omega)}\left(1+\frac{V^{2}}{2(\omega-\Sigma_{\text{f}}(\omega))}\times\right$.

### v2

@@ III.3.2 DC conductivity and resistivity scaling
where $\rho_{c}^{\text{DC}}(T)$ is the resistivity measured precisely at the critical point and $f$ is a universal scaling function. The characteristic energy scale, $T_{0}$, depends on the deviation from the critical interaction, $\delta U=U-U_{c}$, via the power law $T_{0}\sim|\delta U|^{\nu z}$, with $\nu z$ being the dynamical critical exponent. The SBHM exhibits such scaling above the finite temperature critical point, $T_{c}$, with an exponent $\nu z=0.56\pm 0.01$ [terletska2011quantum] when calculated using IPT and CTQMC impurity solvers within DMFT. However, subsequent studies on the same model have reported slightly different values: DMFT+IPT calculations by Vučičević et al. [vuvcivcevic2013finite] yield $\nu z=0.6$, while DMFT+NRG results by Eisenlohr et al. [eisenlohr2019mott] report $\nu z=0.66\pm 0.1$. This variation in the reported exponent is understood to arise from different methodological choices for the Widom line, the trajectory in the temperature-interaction plane used as a reference for the scaling analysis. In this paper, we choose a straight Widom line, similar to the approach discussed in Eisenlohr et al [eisenlohr2019mott].
Since our model undergoes a continuous Mott transition, we expect a similar quantum critical scaling in the electrical resistivity. Here, we perform the scaling analysis at a fixed inter-orbital coupling $t_{\perp}$, while varying the Hubbard interaction $U$. Fig. 9 shows the scaled resistivity $\rho^{\text{DC}}(T)/\rho^{\text{DC}}_{c}(T)$ as a function of the scaled temperature $T/T_{0}$ for different values of $U$. We observe a scaling collapse of the data down to the lowest temperatures, indicating that the scaling persists in the limit $T\rightarrow 0$ on both the metallic and insulating sides. This behavior contrasts sharply with that of the SBHM [terletska2011quantum, vuvcivcevic2013finite], where scaling was observed only above the critical temperature, $T>T_{c}$. An exception is the study by Eisenlohr et al [eisenlohr2019mott], which reported resistivity scaling extending to temperatures below $T_{c}$. However, this scaling at $T<T_{c}$ was limited to the metastable insulating region. In contrast, our results reveal robust resistivity scaling on both sides of the transition as $T\rightarrow 0$, suggesting that the quantum critical point governs the Mott critical transport observed at finite temperatures.
The scaling temperature $T_{0}$ exhibits a power-law dependence, characterized by an exponent of approximately $\nu z_{\text{met}}=0.76$ on the metallic side and $\nu z_{\text{ins}}=0.66$ on the insulating side, as shown in the inset of Fig. 9. The disparity in the exponents across the two sides may originate from the choice of the Widom line, as discussed by Eisenlohr et al. [eisenlohr2019mott], who reported $\nu z\in(0.5,0.8)$ for different Widom line definitions in the SBHM. Our results for the MPAM fall within this range, indicating that the critical scaling of the MPAM is consistent with that of the SBHM.
This finding suggests that the MPAM and SBHM may belong to the same universality class. Consistent with this interpretation, experimental studies on systems exhibiting either first-order or continuous Mott transitions report critical exponents within these bounds. Therefore, the MPAM can be regarded as a prototypical model for exploring Mott quantum criticality.
@@ IV CONCLUSIONS AND OUTLOOK
In this paper, using NRG, we have demonstrated that a modified periodic Anderson model exhibits Mott quantum criticality, characterized by quantum critical scaling in resistivity with a critical exponent comparable to that of the SBHM. This suggests that the MPAM and SBHM may belong to the same universality class. Importantly, the critical temperature $T_{c}$ of the MPAM is zero. We have also shown that the MPAM displays $\omega/T$ scaling, a hallmark of local quantum criticality. Furthermore, we identified isosbestic points in the optical conductivity and showed that the isosbestic points vanish upon approaching the QCP. Analytical expressions for the MIR peak and the charge excitation peak in the optical conductivity were derived in terms of model parameters. We observed that the MIR peak reaches its minimum at the QCP.
A continuous Mott-like metal–insulator transition has been observed in the Falicov–Kimball model [haldar2016quantum, haldar2017quantum], although its critical exponent $\nu z$ is approximately twice that found in the SBHM. Several attempts have been made to tune the finite-temperature critical point of the SBHM to $T=0$ using disorder [aguiar2005effects] or particle–hole asymmetry, but these have been unsuccessful. In contrast, the MPAM features a continuous, symmetry-unbroken transition at zero temperature from a Fermi liquid to a Mott insulating state, while showing a comparable critical exponent $\nu z$ as the SBHM. These findings open up several avenues for future exploration. A central question is whether our results define the universality class for Mott quantum criticality. Another key direction is to explore if a suitable parameter can tune the SBHM to a zero-temperature transition, thereby realizing the physics of the MPAM. Furthermore, investigating the role of disorder in this quantum critical system remains a compelling challenge for future study.
@@ Acknowledgements.
N. S. V. and S. K. K. acknowledge funding from JNCASR and computational resources from the National Supercomputing Mission and JNCASR. S. K. K. acknowledges financial support from the DST-INSPIRE fellowship.
@@ Appendix A Analytic expressions for roots of the Green’s function
$G_{f}$ $(\omega,\epsilon)=\frac{1}{\omega-\Sigma_{\text{f}}(\omega)}\left(1+\frac{V^{2}}{2(\omega-\Sigma_{\text{f}}(\omega))}\times\right$.
@@
Appendix B Derivation of c-band spectral function
In this section, we derive the expression for the c-band spectral function, $\text{A}_{\text{c}}(\omega)$, in the low-frequency limit ($\omega\rightarrow 0$). Consider the Green’s function for the middle conduction band, as given in Eq. 3. Using a partial fraction decomposition, this equation can be rewritten as:
$G_{\text{cc}}$ $(\omega,\epsilon)=\frac{1}{2}\Bigg(\frac{1+\frac{V^{2}}{\sqrt{V^{4}+4(\omega-\Sigma)^{2}\,t_{\perp}^{2}}}}{\omega-\frac{V^{2}}{2(\omega-\Sigma_{\text{f}})}-\sqrt{\frac{V^{4}}{4(\omega-\Sigma_{\text{f}}(\omega))^{2}}+t_{\perp}^{2}}-\epsilon}$
$+\frac{1-\frac{V^{2}}{\sqrt{V^{4}+4(\omega-\Sigma)^{2}\,t_{\perp}^{2}}}}{\omega-\frac{V^{2}}{2(\omega-\Sigma_{\text{f}})}+\sqrt{\frac{V^{4}}{4(\omega-\Sigma_{\text{f}}(\omega))^{2}}+t_{\perp}^{2}}-\epsilon}\Bigg)$ (13)
In the Fermi-liquid regime, we employ a similar argument to that presented in Appendix A to make the low-frequency approximation $(\omega-\Sigma_{\text{f}})\approx\omega/Z$. Substituting this into the expression for $G_{\text{cc}}(\omega,\epsilon)$ yields:
$G_{\text{cc}}(\omega,\epsilon)=\frac{1}{2}\Bigg(\frac{1+\left(1+\frac{4\omega^{2}t_{\perp}^{2}}{Z^{2}V^{4}}\right)^{-\frac{1}{2}}}{\omega-\frac{ZV^{2}}{2\omega}-\sqrt{\frac{Z^{2}V^{4}}{4\omega^{2}}+t_{\perp}^{2}}-\epsilon}$
$\qquad+\frac{1-\left(1+\frac{4\omega^{2}t_{\perp}^{2}}{Z^{2}V^{4}}\right)^{-\frac{1}{2}}}{\omega-\frac{ZV^{2}}{2\omega}+\sqrt{\frac{Z^{2}V^{4}}{4\omega^{2}}+t_{\perp}^{2}}-\epsilon}\Bigg)$ (14)
The c-band spectral function is calculated from the expression $\text{A}_{\text{c}}(\omega)=-\frac{1}{\pi}\text{Im}\Bigg[\int d\epsilon\,\text{A}_{0}(\epsilon)\,G_{\text{cc}}(\omega,\epsilon)\Bigg]$. In the $\omega\rightarrow 0$ limit, simplifying Eq. 14 by using the approximation $(1+x)^{-1/2}\approx 1-x/2,\,\forall\,x\ll 1$. Substituting this result into the integral for $\text{A}_{\text{c}}(\omega)$ yields the final form of the c-band spectral function presented in Subsection III.1.